what are the center and radius of the circle described by the equation (x-5)2+(y+4)2=121?
Answer by Cromlix(4381)

Equation of a circle
(x - a)^2 + (y - b)^2 = r^2
(x-5)2+(y+4)2=121
Centre = (5,-4) radius = 11
Hope this helps.
:-)